I need help please

To a solution containing 0.15 M Ci- ion and 0.15 M Br- ion, you add some solid AgNO3. Ksp for AgCl is 1.6 * 10^-10 and for AgBr is 5.0*10^-13. the addition of solid doesn't change the total volume.
a) which component AgCl or AgBr precipitates first?

b) what is the concentration of the first anion to precipitate when the silver halide of the second anion starts to precipitate?
Answer #1

( a ) solution which requires least concentration of Ag+ or which attain Qsp value first will precipitate first.

[ Ag+ ] require to precipitate AgCl = Ksp ( AgCl ) / [ Cl ] = 1.6 x 10–10 / 0.15 = 1.06 x 10–9 M

[ Ag+ ] requires to precipitate AgBr = Ksp ( AgBr ) / [ Br ] = 5 x 10–13 / 0.15 = 3.33 x 10-12 M

so AgBr will form precipitate first.

( b ) we have to calculate concentration of bromide ion when chloride ion begin to precipitate.

when AgCl form precipitate concentration of silver  ion = 1.06 x 10–9 M

[ Br ] = Ksp ( AgBr ) / [ Ag+ ] = 5 x 10–13 / 1.06 x 10–9 = 4.7 x 10–4 M